OpenStudy (anonymous):

I need help with a calculus question. I know all the steps to get the answer (which i will post) I just dont understand them. So it should be a pretty easy problem to help with

OpenStudy (anonymous):

evaluate the definite integral by the limit definition. the problem is \[\int\limits_{4}^{10} 6 dx\]

OpenStudy (anonymous):

step 1: y=6 on [4,10] (note: \[\Delta x=\frac{ 10-4 }{ n }=\frac{ 6 }{ n } , \left| \left| \Delta \right| \right| \rightarrow 0 (as) n \rightarrow \infty \]

OpenStudy (anonymous):

which part do you need explanation?

OpenStudy (anonymous):

all of it. im not sure of why anything is done. do you want me to finish putting the steps up?

OpenStudy (anonymous):

in order to evaluate the integral, absent formulas to evaluate it, you have to resort to using the limit definition of the integral.

OpenStudy (anonymous):

to solve\[\int_4^{10} 6 dx\] is to evaluate \[\lim_{ n \rightarrow +\infty} \sum_{i=1}^{n}6 \Delta x\]

OpenStudy (anonymous):

\[\sum_{i-1}^{n} F(c _{i}) \Delta x _{i}\] \[\sum_{i-1}^{n} f(4+\frac{ 6i }{ n })(\frac{ 6 }{ n } ) \] \[\sum_{i-1}^{n} 6 (\frac{ 6 }{ n })\] \[\sum_{i-1}^{n}\frac{ 36 }{ n }\] \[\frac{ 1 }{ n }\sum_{i-1}^{n}36\] \[\frac{ 1 }{ n }(36n)\] =36

OpenStudy (anonymous):

the rest of the steps :)

OpenStudy (anonymous):

we state for the record (^^) \(f(x)=6,\,a=4,\,b=10\) then compute the parts of the limit definition.

OpenStudy (anonymous):

\[\Delta x=\frac{6}{n}\]

OpenStudy (anonymous):